OpenStudy (aonz):

Help me please :) (a) Differentiate the semicircle y =sqrt(169 − x^2) , find the equation of the tangent at P(12, 5), and find the x-intercept and y-intercept of the tangent. (b) Show that the perpendicular distance from the tangent to the centre equals the radius. (c) Find the area of the triangle enclosed by the tangent and the two axes. (d) Find the perimeter of this triangle.

OpenStudy (aonz):

i have gotten the gradient to be -12/5

OpenStudy (aonz):

I have done part (a) and equation is 5y+12x-169=0

OpenStudy (anonymous):

First of all, y =169 − x^2 is NOT A SEMICIRCLE it is a PARABOLA...

OpenStudy (aonz):

umm radius is 13? sorry i changed the question

OpenStudy (anonymous):

Okay - that makes more sense!

OpenStudy (anonymous):

Now do you know the centre of the circle?

OpenStudy (aonz):

is it the origin?

OpenStudy (anonymous):

yep, do you know how to get it?

OpenStudy (aonz):

yea ill use the perpendicular formula

OpenStudy (aonz):

now i have gotten the distance to be 13?

OpenStudy (anonymous):

you can do it another way: y = √169-x^2 y^2 = 169 - x^2 y^2 + x^2 = 169 Therefore C (0,0) and r = √169 = 13

OpenStudy (anonymous):

Cool, now for part b) do you know how to find the distance between two points, (0,0) and (12,5) ?

OpenStudy (anonymous):

@AonZ @BarbaraKara Can I be of any assistance?

OpenStudy (aonz):

sure :)

OpenStudy (anonymous):

What do you need help with atm @AonZ

OpenStudy (aonz):

part (c) so far

OpenStudy (aonz):

it should be ok, i think i can solve this now

OpenStudy (anonymous):

So the equation of the tangent line is 5y + 12x - 169 = 0. Is that correct? @AonZ

OpenStudy (aonz):

I think im fine now, thanks :)

OpenStudy (anonymous):

well ok. tag me if you need any help.

OpenStudy (aonz):

Sure, will do